Operations Report – November 29th

MDRS Operations Report 27-29-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: One of the charge controllers appears to have frozen on ABSORB setting. I noticed it yesterday and confirmed it today.

Action taken for non-nominal systems: I was not able to reset it so I will equalize it tomorrow. I didn’t want to do anything to change the system until I tried to autostart one more time and equalizing it requires that charge controller to shut down and restart.

Generator: The power system is giving me no love. For three nights I’ve set the generator to autostart and it has not. I’m giving it one more chance before I give up and run it manually. The good news is that now that the propane installation had been corrected, I can do that if I need to. I reprogrammed everything back to use SOC as the parameter for the autostart and let’s see what happens tonight. If this doesn’t work, I’m going with Occam’s Razer and assuming the problem is in the power system’s wiring, as this is the second generator that won’t autostart. But we would need to wait until spring to resolve that as our installer is out of the country. I’ve been fighting this how many years now?

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours:

Average 51.37

Minimum 46.50

Maximum 58.80

Since the online system is not showing SOC, I am assuming my changes to the power system did not work. Should I just bang my head against a wall now?

Notes on solar: Nothing to report

Propane Reading, station tank – 46 %

Propane Reading, director tank— 28 %

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 300

Static to Loft Pump used – yes

Water in GreenHab – 20 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 60 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town and to quarry with media team

CrewCar used and why, where? No

Luna used and why, where? No

General notes and comments: Propane company should be here tomorrow to switch out the three Amerigas tanks.

Summary of internet: All accounts are nominal

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Turned on the main heater all day yesterday. No evidence of a propane leak.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: I’ve been working in the GreenHab getting it ready for the next crew. Everything is growing well. I’ve turned on the old lights for the past three nights. They have not run as programmed. I changed the timer today to an older one and we will see what happens. We may need to order a new one.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: It’s good to see the dome open when I go out right before bed and first thing in the morning.

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Atila and Sebastian will be arriving tomorrow so this will be the last night I spend alone!

Operations Report Nov 25th – 26th

MDRS Operations Report 25-26-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: I worked on the power system yesterday and today. I was getting a little desperate! Late this afternoon, it appears we are currently operational with the generator in autostart using VDC the parameter. Keep your fingers crossed! I did not change any of the power usage to make sure this actually works before heating the Hab and turning on the supplemental lighting.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Friday):

Average 51.68

Minimum 47.18

Maximum 58.58

SOC Last 24 hours:

100 Max

68 Min

86.5 Average

Notes on solar: Nothing to report

Propane Reading, station tank – 50 %

Propane Reading, director tank— 36%

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 340

Static to Loft Pump used – yes

Water in GreenHab – 20 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? No

Luna used and why, where? Yes, to town

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Nothing to report. Still not heating the building and the water is freezing lightly at night, then thawing during the day.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Nothing to report except there are too many greens ready for harvest for just one person.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report Nov 21st – 24th

MDRS Operations Report 21-24-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Propane line has been changed to ¾ inch per manufacturer specs. In addition, since we calculated a possible use of 20 gallons a night in the middle of winter, we exchanged the 250-gallon tank for a 500-gallon tank.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Wednesday):

Average 51.20

Minimum 46.40

Maximum 58.40

Notes on solar: Tomorrow I will reprogram the generator autostart back into the power system.

Propane Reading, station tank – 50 %

Propane Reading, director tank— 40%

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 370 gallons (I plugged in the heat tape as the Hab’s water lines have begun to freeze at night)

Static to Loft Pump used – yes

Water in GreenHab – 20 gallons (not accessible)

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): Nothing to report

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? Yes, to start. Started but was sluggish. Will replace battery before crew use.

Luna used and why, where? Yes, to town

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: A survey of water line issues show that all three are simple fixes and will be completed soon. The toilet water line is closed and the new part arrived today. The water lines are freezing during the night and thawing by afternoon.

Summary of Outpost operations: In my trailer, a new kitchen faucet and a new bathroom were installed. The shower was fixed to operate minimally. Two 250-gallon propane tanks have been dropped off to be switched over for both trailers after Thanksgiving.

Summary of GreenHab operations: Using about 10 gallons of water a day for the plants currently growing. The LED lights arrived today and need to be installed. I’m keeping the temperature at night around 77 degrees F.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report – November 20th

MDRS Operations Report 17-20-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Propane line is being changed from ½ inch to ¾ inch as required by the manufacturer’s specs. This was started Friday and will be completed Monday as the installer needed to get more parts.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Saturday):

Average 51.40Minimum 46.50
Maximum 58.40

Notes on solar: nothing to report

Propane Reading, station tank – 50 %
Propane Reading, director tank— 40%
Propane Reading, intern tank— 72+ %
Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 30 gallons
Water Meter: not checked units
Water (static tank) – 400 gallons

Static to Loft Pump used – yes

Water in GreenHab – 20 gallons (not accessible)
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: yes

Perseverance rover used: no

Hours:

Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: yes

Hours:

Beginning charge: 100
Ending charge: 89

Currently charging: no

Spirit rover used: no

Hours:

Beginning charge: 100
Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100
Ending charge: 100

Currently charging: yes

Curiosity rover used: no
Hours:

Beginning charge: 100
Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): The 300 was returned to the station today, completely inoperable. 350.3 was removed from the station in exchange

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Currently only 350.1 is operational. 350.2 runs, but the accelerator sticks, Honda still awaiting two rear tires.

HabCar used and why, where? Yes, to town

CrewCar used and why, where? No

Luna used and why, where? No

General notes and comments: Updated the tags and paperwork on the CrewCar. It started today with a drag on the battery.

Summary of internet: All accounts are nominal.

Summary of suits and radios: Nothing to report.

Campus-wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Did another propane check Friday, no evidence of a leak. I did smell something when I entered the Hab, which I believe to be sewer gas. I did another check Saturday, no smell and no evidence of a leak.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: I’m harvesting some plants as I’m the only one here to eat them. I am using the water from the static tank to water the plants. Yes, I am so lazy that I’m waiting for Atila to arrive so he can haul water. In my defense, he likes it and I don’t. I’ve ordered some LED grow lights that use considerably less power than what we currently have.
Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Assistant Director will be back on campus tomorrow!

Operations Report – November 13-16th

MDRS Operations Report 13-16-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: nothing to report

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Tuesday):

Average 50.85

Minimum 47.50

Maximum 58.0

Notes on solar: Cloudy until this afternoon but SOC was 99% at sunset

Propane Reading, station tank – not checked %

Propane Reading, director tank— not checked %

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): not checked gallons

Water Meter: not checked units

Water (static tank) – not checked gallons

Static to Loft Pump used – no

Water in GreenHab – 20 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: no

Notes on rovers: Unplugged due to the cloudy weather today

ATV’s Used: (Honda, 350.1, 350.2, 350.3): no

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? Yes, to town to dump trash

CrewCar used and why, where? No

Luna used and why, where? Yes, to Bicknell

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal.

Summary of suits and radios: Nothing to report.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: nothing to report

Summary of Hab operations: Did a complete propane check on Monday and still no evidence of any propane problems. Heaters are off, but the hot water heater, which was what the crew reported as leaking is working nominally.

Summary of Outpost operations: My shower stopped working. No idea why.

Summary of GreenHab operations: Fertilized the plants. Between this and the increase in temperature, I have seen an increase in growth. I need to get the supplemental lighting set up but worry about the increase in nighttime power use. I am looking into switching what we currently have with LED lights.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at report.

Operations Report Nov 11th – 12th

MDRS Operations Report 11-12-NOV-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: n/a

Generator: Propane installer confirmed the set up does not meet the requirements of the generator and will be here next week to redo the work.

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— VDC Last 24 hours (Friday):

Average 52.04

Minimum 48.30

Maximum 58.0

Notes on solar: Nothing to report

Propane Reading, station tank – 64%

Propane Reading, director tank— 55%

Propane Reading, intern tank— 72+ %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): not checked gallons

Water Meter: not checked units

Water (static tank) – not checked gallons

Static to Loft Pump used – no

Water in GreenHab – 50 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 80 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Not sure if this was reported, but Percy’s batteries were refilled during SWB.

ATV’s Used: (Honda, 350.1, 350.2, 350.3): no

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: I have scheduled servicing for all four vehicles.

HabCar used and why, where? Yes. I drove it to the Hab to replace the bent that was attached to the pump in the water tank. I accidently bumped into the front porch railing and put a small dent in the liftgate.

CrewCar used and why, where? Yes, to town for packages

Luna used and why, where? No

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal.

Summary of suits and radios: Nothing to report.

Campus wide inspection, if action taken, what and why? The problem with the foundation of the ScienceDome is getting worse. We need to build up the sand base around it and add gravel to keep it in place or water will continue to undermine the structure.

Summary of general operations: The septic was pumped out on Thursday.

Summary of Hab operations: I have checked for a propane leak two more times yesterday and once today. I have used both meters, the alarm and soapy water. In addition, I turned on and off all appliances to see if I could create a leak. There is still no evidence that there has or is a propane leak in the Hab. I have scheduled repair of all plumbing issues and am ordering parts for replacement/repair.

Summary of Outpost operations: I have scheduled repair of all plumbing issues and am ordering parts for replacement/repair.

Summary of GreenHab operations: Watered plants. Increased nighttime temperature back to 75 degrees F. Added two smaller plastic watering cans.

Summary of ScienceDome operations: Watered plants

Summary of RAM operations: Added to tool inventory: 5 new Phillips screwdrivers, two eyeglass screwdriver sets, more bits for the cordless drill.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing at this time.

Crew 239 Journalist Report for November 10th

Journalist Report 10 Nov 2021

By Emily Milne

MDRS Crew 239 Sol 3 (11/10/2021):

We began our morning with a breakfast of oatmeal, Cliff bars, tea, and coffee before our 1000 EVA. The EVA was a return to rover training by Kissing Camel Ridge and consisted of team leader GHO Cynthia Montanez, HSO Emily Milne, and ENG Aidan Moga. ENG Moga took samples for analysis in his spectrometer to find potential landing sites. Upon returning at 1202 CO Crisman made jambalaya for Lunch, GHO Montanez transplanted and watered the plants, ENG Moga utilized the spectrometer, and the rest of the crew worked on individual research. HSO Milne cooked pasta with gravy for dinner using tomato powder, tomato chunks, onion, and pork sausage. After eating, the crew watched “The Thing” before retiring for the night.

Crew 239 Operations Report for November 10th

Crew 239 Operations Report 10-11-2021

Name of person filing report: Aidan Moga

Non-nominal systems: Toilet, Propane supply to generator, water heater

Action taken for non-nominal systems: The leak in the toilet fitting has gotten worse. The crew is working on better diagnosing and solving the problem later tonight. The power supply situation was better today, due to the sunny weather. The crew was able to charge the rovers and devices during the daytime. Power rationing continues in the afternoon/evenings. At around 18:20 the water heater was on and started to leak gas into the lower level of the hab. The crew attempted to use the gas detector from the ram to locate the leak. However, the leak couldn’t be replicated and wasn’t precisely located. The crew will continue to monitor the situation.

Daily crew water use: 21 gallons

GreenHab water use: 3 gallons

Hab toilet tank emptied: No

Rovers used: Yes, Spirit and Opportunity.

Notes on rovers (report nominal or any anomalies): Nominal

Summary of crew and Mission Support internet: Nominal

Summary of suits and radios: During EVA pause, the fan on GHO’s suit 7 sounded like it was being interrupted. The suit was set aside and replaced with suit 8 for the EVA. When suit 7 was tested later, the fan appeared to operate nominally. For now, suit 8 will continue to be used. Some of the radios have damaged antennas; these have been put aside to not be used during EVAs.

Campus wide inspection, if action taken, what and why? Nominal

Summary of general operations: Nominal

Summary of Hab operations: Nominal

Summary of GreenHab operations (note any harvested crops):

· Temperature throughout Sol 2:

o 1:00 p.m. -> 110 degree F with 7% humidity

o 4:34 p.m. -> 72.5 degree F with 19% humidity

o 6:30 p.m. -> 70.6 degree F with 19 % humidity

· Hours of Supplemental Light: None

· Time(s) of watering for crops:

o At 1:19 p.m. I used 1 gallon

o At 4:54 p.m. I used 3 gallons

· Narrative:

o Transplanted Thai pepper plant to a larger pot

o Transplanted largest cucumber plant to a larger pot and straightened the stems.

o Moved smaller cucumbers to a location furthest away from the door to help with climate control.

o Transplanted two large sweetie tomatoes into seven smaller pots

o Transplanted regular, but undefined, tomato plant to a larger pot.

o Discovered that the zucchini’s do not have a disease, instead they are ‘burning’ due to the extreme temperatures the GreenHab has. I will handle accordingly to prevent any more damage on these plants.

· Support/ Supplied Needed:

o Requesting more potting soil

o Requesting red wagon within the GreenHab to be removed.

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns, supplies needed and requests: None